# A Motivic Pairing and the Mellin Transform in Function Fields

### v2

@@ 1. Introduction
We now briefly state a special case of the main theorem of this paper and explain why this should be viewed as the correct analogue of the Mellin transform. Let $q=p^{r}$ be a prime power, and let $A=\mathbb{F}_{q}[\theta]$ and $K=\mathbb{F}_{q}(\theta)$ (in most sections of the paper we actually prove our results for an arbitrary function field). Let $K_{\infty}=\mathbb{F}_{q}((1/\theta))$ and let $\mathbb{C}_{\infty}$ be a completion of an algebraic closure of $K_{\infty}$. Set $u=\frac{\tilde{\pi}}{t-\theta}\in\mathbb{C}_{\infty}(t)$, where $\tilde{\pi}$ is the Carlitz period (see (5.5)) and $t$ is an independent variable. For a specified element $\mathbf{z}\in\mathbb{C}_{\infty}$ (here $n$ is the dimension of a particular $t$-module) we define a map $\delta_{1,\mathbf{z}}^{M}$ from a $t$-motive $M$ to $\mathbb{C}_{\infty}$, which we view as an analogue of integration. We then have the following special case of our main theorem.
@@
$\delta^{M}_{1,\mathbf{z}}\left(-\frac{u}{\exp_{C}(u)}\right)=\Gamma_{A}(n)\zeta_{A}(n)\in K_{\infty},$
@@
The main results of Sections 3 and 4 give explicit formulas for the coefficients of the exponential and logarithm functions of an arbitrary Anderson $A$-module subject to a couple technical conditions (see Defs. 2.3 and 2.12). These formulas are contained in Corollaries 3.9 and 4.5. Our main technique is to construct two pairings between the associated $A$-motive and dual $A$-motive of $\phi$ (see Def. 2.3 for the definition of $A$-motives). We then show that a specialization of these parings returns the exponential and logarithmic functions and this allows us to prove precise formulas for $\operatorname{Exp}_{\phi}$ and $\operatorname{Log}_{\phi}$ in Theorems 3.8 and 4.4. We view this $A$-motivic set up as being a very natural setting to study these coefficients, and our formulas subsume many previous ad hoc approaches (see [21], [14], [9]). As an immediate application of our formulas, we give a sufficient condition for the exponential and logarithm coefficients of $t$-modules to be invertible in Theorems 4.7 and 4.8.
@@ 2. Review of Anderson Motives and Modules
In this section we review the definitions and much of the theory involved in $A$-motives, dual $A$-motives and Anderson $A$-modules. Most of these results are originally due to Anderson [1], although here we follow the exposition of Hartl and Juschka [18]. We also prove several new results which will be used in subsequent sections to define our motivic pairings. Let $\mathbb{F}_{q}$ be the finite field of size $q=p^{r}$ and let $X$ be a projective, smooth, geometically connected curve defined over $\mathbb{F}_{q}$ with fixed point at infinity $\infty\in X(\mathbb{F}_{q})$. Let $K$ be the function field of $X$ and let $A$ be the functions on $X$ regular away from $\infty$. Let $K_{\infty}$ be the completion of $K$ at the place of $K$ corresponding to $\infty$, and let $\mathbb{C}_{\infty}$ be the completion of an algebraic closure of $K_{\infty}$. Let $L$ be an algebraically closed subfield with $K\subset L\subset\mathbb{C}_{\infty}$. We comment that in some subsections we will set $L=\mathbb{C}_{\infty}$, and the reader doesn’t lose much generality by simply making this assumption throughout the paper. We view $L\otimes_{\mathbb{F}_{q}}A$ as the ring of functions regular away from the point $\infty$ of the curve $L\times_{\mathbb{F}_{q}}X$ and view $\operatorname{Frac}(L\otimes_{\mathbb{F}_{q}}A)$ as its function field. We fix an inclusion $\mathbb{F}_{q}(t)\subset K$ and let $\Xi\in X(K)$ be a pull back of the point $t\in\mathbb{P}^{1}(\mathbb{F}_{q}(t))$ under the corresponding projection $X\to\mathbb{P}^{1}$.
@@
    As an $\mathbb{F}_{q}$-vector space scheme over $L$, the scheme $E$ is isomorphic to $\mathbb{G}_{a}^{d}$
@@
We define a ring homomorphism $*:(L\otimes A)[\tau]\to(L\otimes A)[\sigma]$, for $\alpha=\sum_{i=0}^{m}a_{i}\tau^{i}\in(L\otimes A)[\tau]$, by setting
@@
Let $M_{K}=M\otimes_{L\otimes_{\mathbb{F}_{q}}A}\operatorname{Frac}(L\otimes_{\mathbb{F}_{q}}A)$ be the $(\operatorname{Frac}(L\otimes_{\mathbb{F}_{q}}A))[\tau]$-module where the $\tau$-action on $M$ is as given above and the $\tau$-action on $\operatorname{Frac}(L\otimes_{\mathbb{F}_{q}}A)$ acts as the Frobenius (as in Def. 2.6). Define $N_{K}=N\otimes_{L\otimes_{\mathbb{F}_{q}}A}\operatorname{Frac}(L\otimes_{\mathbb{F}_{q}}A)$ similarly where the $\sigma$ acts on $\operatorname{Frac}(L\otimes_{\mathbb{F}_{q}}A)$ as the inverse of Frobenius. Note that $M_{K}$ and $N_{K}$ are $\operatorname{Frac}(L\otimes_{\mathbb{F}_{q}}A)$-free modules of rank $r$.
@@
We comment that the condition in (1) of the above theorem that the $\tau$- and $\sigma$-action be represented by matrices which are transposes of each other is not required for the main theorems in this paper — although it is convenient for calculations. The condition on the $b$-action in part (2) is required for our main theorems.
Definition 2.12.
We fix the bases of Lemma 2.10 for use throughout the paper. Further, given such bases and the matrix $\Phi$ as defined above, we say that the Anderson $A$-module $\phi$ is $\Xi$-regular if $\det(\Phi)$ and each coordinate of $\Phi$ is regular at $\Xi^{(i)}$ for all $i\in\mathbb{Z}$. Note that if $A=\mathbb{F}_{q}[\theta]$, then $\Phi\in\operatorname{Mat}_{d}(\mathbb{C}_{\infty}[t])$ and $\det(\Phi)\in(t-\theta)^{d}\mathbb{C}_{\infty}[t]$, thus all $t$-modules are automatically $\Xi$-regular. Throughout the rest of the paper we will assume that all our Anderson $A$-modules are $\Xi$-regular.
@@
then writing (notice the reversal of order in the vector compared Definition 2.13)
@@
Thus, both $\delta_{1}^{M}$ and $\delta_{1}^{N}$ allow us to recover the Anderson $A$-module $\phi$, while both $\delta_{0}^{M}$ and $\delta_{0}^{N}$ allow us to recover the Lie action of $\phi$.
@@
For $c\in\mathbb{C}_{\infty}$, we define the Tate algebra in the variable $t$ with radius $c$
@@
Note that this ring is a ring of regular functions on an open rigid analytic subspace of the rigid analytic space associated to our curve $X$. We refer the reader to [18, §2.3.3] for full details on the rigid analytic geometry. We let $\tau$ (and $\sigma$) act on $\mathbb{Y}_{c}$ diagonally via Frobenius twisting (inverse twisting). Thus we let $L=\mathbb{C}_{\infty}$ and set $M_{\theta}=M\otimes_{(\mathbb{C}_{\infty}\otimes A)}\mathbb{Y}_{\theta}$ and $N_{\theta}=N\otimes_{(\mathbb{C}_{\infty}\otimes A)}\mathbb{Y}_{\theta}$ with diagonal $\tau$ and $\sigma$ action, respectively. We set $N_{\Gamma}$ to be the submodule of $N_{\theta}$ consisting of all $h\in N_{\theta}$ such that $\sigma^{i}(h)\in N_{\theta}$ for all $i\in\mathbb{Z}$.
@@
$\lVert h\rVert_{c}=\sup|c|_{\infty}^{i}\cdot|b_{i}|_{\infty},$
where $|\cdot|_{\infty}$ is the norm used for the completion $K_{\infty}$. We then extend this norm to $\mathbb{Y}_{c}$ by recognizing that $\mathbb{Y}_{c}\cong\mathbb{T}_{c}^{\ell}$ where $\ell$ is the degree of $K/\mathbb{F}_{q}(t)$, and taking the max of the norm in each coordinate. Finally, we extend this norm to matrices $\operatorname{Mat}_{m\times m}(\mathbb{Y}_{c})$, again by taking the max of the norm of each coordinate. Note that this matrix norm is sub-multiplicative, so for $B_{1},B_{2}\in\operatorname{Mat}_{m\times m}(\mathbb{Y}_{c})$ we have
@@
The proof of the first extension of $\delta_{0}^{N}$ is given in [18, Prop 2.5.8] and relies on the fact that the map $\delta_{0}^{N}$ factors through $N/J^{d}N$ ($J$ is the maximal ideal generated by $(1\otimes a)-(a\otimes 1)$). For the second, we note that $M/\tau M$ is a $J$-torsion $L\otimes A$-module, so $\delta_{0}^{M}M\to M/\tau M\cong\operatorname{Lie}_{\phi}(L)$ factors uniquely as
@@
We let $N\langle\sigma\rangle$ denote the subset of $n\in N\otimes_{L[\sigma]}L[[\sigma]]$ such that we have $(c_{1,j},\dots,c_{d,j})^{(j)}\to 0$ as $j\to\infty$ in the above expansion for $n$. Thus there exists an extension of the map $\delta_{1}^{N}:N\langle\sigma\rangle\to\mathbb{C}_{\infty}^{d}$ defined by applying $\delta_{1}^{N}$ to partial sums of (2.9), then taking the limit. For the $A$-motive $M$ and for fixed $\mathbf{z}\in L^{d}$, we define $M\langle\tau\rangle$ similarly, except subject to the condition that $(c_{1,j},\dots,c_{d,j})\mathbf{z}^{(j)}\to 0$. As above, there exists an extension of $\delta_{M,\mathbf{z}}^{1}$ to $M\langle\tau\rangle$.
@@ 3. The Exponential Motivic Pairing
The goal of this section is to define a pairing between the rings $(L\otimes_{\mathbb{F}_{q}}A)[\tau]$ and $(L\otimes_{\mathbb{F}_{q}}A)[\sigma]$, as well as a related pairing between the motives $M$ and $N$, which specializes to give the exponential function. This will allow us to give extremely explicit formulas for the coefficients of the exponential function for any abelian $\Xi$-finite Anderson $A$-module.
@@
We recall for the reader’s convenience that $\delta_{0}^{M}(g)\in L^{d}$ for any $g\in M_{(J)}$. Thus, in the above summand, $\delta_{0}^{M}(\tau^{-i}(x(g_{k})))^{\top}\mathbf{z}\in L$, and thus it makes sense to multiply this by $\sigma^{i}(y(h_{k}))\in N$. This is an element of $N$, so we can evaluate it under $\delta_{1}^{N}$, and taking the sum for $i\geq 0$ gives an element $\operatorname{Mat}_{d}(\mathbb{C}_{\infty})[[\mathbf{z}]]$. The question remains whether this sum converges to an element in $\mathbb{C}_{\infty}^{d}$. For now, we view the pairing as a formal sum in $\operatorname{Mat}_{d}(\mathbb{C}_{\infty})[[\mathbf{z}]]$, and we will address the convergence issues in Lemma 3.11 after establishing some of the properties of $F$.
@@
We thank the referee for the following helpful suggestion. It is possible to give a completely coordinate free definition of (3.2) (without the deformation parameters $x,y$). Namely, for $m\in M$ and $n\in N$ and $\mathbf{z}\in\operatorname{Lie}_{E}(L)$ we set
@@
where we comment that this definition makes sense because we identify $\operatorname{Lie}_{E}(L)\cong M/\tau M\cong\operatorname{Hom}_{L}(\operatorname{Lie}_{E}(L),L)$ as in [18, 2.5.2], thus $\delta_{0}^{M}(\tau^{-i}m)(z)$ makes sense and is in $L$. Then, $N$ is an $L$-vector space, so $\delta_{0}^{M}(\tau^{-i}m)(z)\cdot\sigma^{i}n\in N$. Finally, $\delta_{1}^{N}\left(\delta_{0}^{M}(\tau^{-i}m)(z)\cdot\sigma^{i}n\right)\in N/(1-\sigma)N\cong E(L)$.
Then, to recover Definition 3.1 we evaluate the above function $F$ at a chosen basis for $M$ and its dual basis for $N$ and take the trace
@@
For $\mathbf{z}\in L^{d}$, as formal series in $\operatorname{Mat}_{d}(\mathbb{C}_{\infty})[[\mathbf{z}]]$ we have
@@
Since $F(1,1;\mathbf{z})$ and $\operatorname{Exp}_{\phi}(\mathbf{z})$ agree as formal power series, and since we know $\operatorname{Exp}_{\phi}(\mathbf{z})$ is an entire function on $\mathbb{C}_{\infty}^{d}$ (see [1]), we conclude that $F(1,1;\mathbf{z})$ is also an entire function of $\mathbf{z}$. By Proposition 3.6 we see that $F(cx,dx;\mathbf{z})=cdF(x,y;\mathbf{z})$ for $c,d\in L$, that $F(ax,bx;\mathbf{z})=\phi_{b}\circ F(x,y;d[a]\mathbf{z})$ for $a,b\in A$. Further, recall that for $z\in\mathbb{C}_{\infty}$ and $h_{k}$ a basis element of Lemma 2.10, if $\delta_{1}^{N}(z\sigma^{i}(h_{k}))=w\in\mathbb{C}_{\infty}^{d}$, then by (2.4) we have that $\delta_{1}^{N}(z\sigma^{i+1}(h_{k}))=w^{(1)}$. From this, it follows that $F(\tau,1;\mathbf{z})=F(1,\sigma;\mathbf{z})=F(1,1;\mathbf{z})^{(1)}$. Thus, the convergence of $F(x,y;\mathbf{z})$ reduces to the convergence of $F(1,1;\mathbf{z})$, which is an entire function. Thus $\sum_{i=0}^{\infty}\sum_{k=1}^{d}\delta_{0}^{M}(\tau^{-i}(x(g_{k})))^{\top}\mathbf{z}\sigma^{i}(y(h_{k})))\in N\langle\sigma\rangle$.
@@
In this section we give a detailed example of many of the above constructions in the case of the $n$th tensor power of the Carlitz module for fixed $n>0$. We have $K=\mathbb{F}_{q}(t)$ and $A=\mathbb{F}_{q}[t]$ and we set $L=\mathbb{C}_{\infty}$. Let us identify $1\otimes A=\mathbb{F}_{q}[t]$ and $A\otimes 1=\mathbb{F}_{q}[\theta]$ for independent variables $t$ and $\theta$, thus obviating the need for the tensor product in this situation (thus $\mathbb{C}_{\infty}$ is written using the $\theta$ variable). We may realize the $t$-motive $M$ and the dual $t$-motive $N$ as $M=N=\mathbb{C}_{\infty}[t]$ (here $L\otimes_{\mathbb{F}_{q}}A=\mathbb{C}_{\infty}\otimes_{\mathbb{F}_{q}}A=\mathbb{C}_{\infty}[t]$) with
@@
for $z\in M_{K}$. Finally, Corollary 3.9 (here $d=1$) shows that the exponential coefficients $P_{i}$ are given by
$P_{i}=\delta_{0}^{M}(\tau^{-i}(1))=\left(\frac{1}{f^{(-1)}\cdots f^{(-i)}}\bigg{|}_{\Xi}\right)^{(i)}=\frac{1}{f\cdots f^{(i-1)}}\bigg{|}_{\Xi^{(i)}},$

### v3

@@ 1. Introduction
We now briefly state a special case of the main theorem of this paper and explain why this should be viewed as the correct analogue of the Mellin transform. Let $q=p^{r}$ be a prime power, and let $A=\mathbb{F}_{q}[\theta]$ and $K=\mathbb{F}_{q}(\theta)$ (in most sections of the paper we actually prove our results for an arbitrary function field). Let $K_{\infty}=\mathbb{F}_{q}((1/\theta))$ and let $\mathbb{C}_{\infty}$ be a completion of an algebraic closure of $K_{\infty}$. Set $u=\frac{\tilde{\pi}}{\theta-t}\in\mathbb{C}_{\infty}(t)$, where $\tilde{\pi}$ is the Carlitz period (see (5.5)) and $t$ is an independent variable. For a specified element $\mathbf{z}\in\mathbb{C}_{\infty}$ (here $n$ is the dimension of a particular $t$-module) we define a map $\delta_{1,\mathbf{z}}^{M}$ from a $t$-motive $M$ to $\mathbb{C}_{\infty}$, which we view as an analogue of integration. We then have the following special case of our main theorem.
@@
$\delta^{M}_{1,\mathbf{z}}\left(\frac{u}{\exp_{C}(u)}\right)=\Gamma_{A}(n)\zeta_{A}(n)\in K_{\infty},$
@@
The main results of Sections 3 and 4 give explicit formulas for the coefficients of the exponential and logarithm functions of an arbitrary Anderson $A$-module subject to a couple technical conditions (see Defs. 2.3 and 2.11). These formulas are contained in Corollaries 3.9 and 4.5. Our main technique is to construct two pairings between the associated $A$-motive and dual $A$-motive of $\phi$ (see Def. 2.3 for the definition of $A$-motives). We then show that a specialization of these parings returns the exponential and logarithmic functions and this allows us to prove precise formulas for $\operatorname{Exp}_{\phi}$ and $\operatorname{Log}_{\phi}$ in Theorems 3.8 and 4.4. We view this $A$-motivic set up as being a very natural setting to study these coefficients, and our formulas subsume many previous ad hoc approaches (see [21], [14], [9]). As an immediate application of our formulas, we give a sufficient condition for the exponential and logarithm coefficients of $t$-modules to be invertible in Theorems 4.7 and 4.8.
@@ 2. Review of Anderson Motives and Modules
In this section we review the definitions and much of the theory involved in $A$-motives, dual $A$-motives and Anderson $A$-modules. Most of these results are originally due to Anderson [1], although here we follow the exposition of Hartl and Juschka [18]. We also prove several new results which will be used in subsequent sections to define our motivic pairings. Let $\mathbb{F}_{q}$ be the finite field of size $q=p^{r}$ and let $X$ be a projective, smooth, geometically connected curve defined over $\mathbb{F}_{q}$ with fixed point at infinity $\infty\in X(\mathbb{F}_{q})$. Let $K$ be the function field of $X$ and let $A$ be the functions on $X$ regular away from $\infty$. Let $K_{\infty}$ be the completion of $K$ at the place of $K$ corresponding to $\infty$, and let $\mathbb{C}_{\infty}$ be the completion of an algebraic closure of $K_{\infty}$. Let $L$ be an algebraically closed subfield with $K\subset L\subset\mathbb{C}_{\infty}$. We comment that in some subsections we will set $L=\mathbb{C}_{\infty}$, and the reader doesn’t lose much generality by simply making this assumption throughout the paper. We view $L\otimes_{\mathbb{F}_{q}}A$ as the ring of functions regular away from the point $\infty$ of the curve $L\times_{\mathbb{F}_{q}}X$ and view $\operatorname{Frac}(L\otimes_{\mathbb{F}_{q}}A)$ as its function field. Finally, we let $\Xi$ be the point of $L\times_{\mathbb{F}_{q}}X$ corresponding to the kernel of the multiplication map, $L\otimes_{\mathbb{F}_{q}}A\to L$ given by $\sum l_{i}\otimes a_{i}\mapsto\sum l_{i}a_{i}$.
@@
    As an $\mathbb{F}_{q}$-vector space scheme over $L$, we have $E$ is isomorphic to $\mathbb{G}_{a}^{d}$.
@@
We define a ring homomorphism $*:(L\otimes A)[\tau]\to(L\otimes A)[\sigma]$ by setting, for $\alpha=\sum_{i=0}^{m}a_{i}\tau^{i}\in(L\otimes A)[\tau]$,
@@
Let $M_{K}=M\otimes_{L\otimes_{\mathbb{F}_{q}}A}\operatorname{Frac}(L\otimes_{\mathbb{F}_{q}}A)$ be the $(\operatorname{Frac}(L\otimes_{\mathbb{F}_{q}}A))[\tau]$-module where the $\tau$-action on $M$ is as given above and the $\tau$-action on $\operatorname{Frac}(L\otimes_{\mathbb{F}_{q}}A)$ acts as the Frobenius (as in Def. 2.6). Define $N_{K}=N\otimes_{L\otimes_{\mathbb{F}_{q}}A}\operatorname{Frac}(L\otimes_{\mathbb{F}_{q}}A)$ similarly where the $\sigma$ acts on $\operatorname{Frac}(L\otimes_{\mathbb{F}_{q}}A)$ as the inverse of Frobenius. Note that $M_{K}$ and $N_{K}$ are $\operatorname{Frac}(L\otimes_{\mathbb{F}_{q}}A)$-vector spaces of dimension $r$.
@@
Definition 2.11.
We fix the bases of Lemma 2.10 for use throughout the paper. Further, given such bases and the matrix $\Phi$ as defined above, we say that the Anderson $A$-module $\phi$ is $\Xi$-regular if each coordinate of $\Phi$ is regular at $\Xi^{(i)}$ for all $i\in\mathbb{Z}$. Note that if $A=\mathbb{F}_{q}[\theta]$, then $\Phi\in\operatorname{Mat}_{d}(L[t])$ and $\det(\Phi)\in(t-\theta)^{d}L[t]$, thus all $t$-modules are automatically $\Xi$-regular. Throughout the rest of the paper we will assume that all our Anderson $A$-modules are $\Xi$-regular.
@@
We also comment here, thanks to excellent suggestions by the referee, that many of the constructions in this paper can be done without assuming $\ \Xi$-regularity. Notably the construction of the pairings $F$ and $G$ of Sections 3 and 4 can be done without it, as is explained in Remark 3.3. In fact, it is a widely assumed folklore conjecture that all Anderson $A$-modules are $\Xi$-regular, so we hope that with future work this assumption can be removed entirely.
@@
then writing
@@
We comment that $\delta_{0}^{M}$ and $\delta_{0}^{N}$ can be recovered by the maps induced by $M\to M/\tau M$ and $N\to N/\sigma N$. Similarly, $\delta_{1}^{N}$ can be recovered by the map induced by $N\to N/(1-\sigma)N$, and while it is true that $\delta_{1,\mathbf{z}}^{M}$ can be recovered by the map induced by $N\to N/(1-\tau)N$, this is much more subtle.
@@
For $c\in\mathbb{C}_{\infty}$, we define the Tate algebra in the variable $t$ with radius $|c|_{\infty}$, where $|\cdot|_{\infty}$ is the norm used for the completion $K_{\infty}$.
@@
Note that this ring is a ring of regular functions on an open rigid analytic subspace of the rigid analytic space associated to our curve $X$. We refer the reader to [18, §2.3.3] for full details on the rigid analytic geometry (in the case $d_{\infty}=1$). We let $\tau$ (and $\sigma$) act on $\mathbb{Y}_{c}$ diagonally via Frobenius twisting (inverse twisting). Thus we let $L=\mathbb{C}_{\infty}$ and set $M_{\theta}=M\otimes_{(\mathbb{C}_{\infty}\otimes A)}\mathbb{Y}_{\theta}$ and $N_{\theta}=N\otimes_{(\mathbb{C}_{\infty}\otimes A)}\mathbb{Y}_{\theta}$ with diagonal $\tau$ and $\sigma$ action, respectively. We set $N_{\Gamma}$ to be the submodule of $N_{\theta}$ consisting of all $h\in N_{\theta}$ such that $\sigma^{i}(h)\in N_{\theta}$ for all $i\in\mathbb{Z}$.
@@
$\lVert h\rVert_{c}=\sup|c|_{\infty}^{i}\cdot|b_{i}|_{\infty}.$
We then extend this norm to $\mathbb{Y}_{c}$ by recognizing that $\mathbb{Y}_{c}\cong\mathbb{T}_{c}^{\ell}$ where $\ell$ is the degree of $K/\mathbb{F}_{q}(t)$, and taking the max of the norm in each coordinate. Finally, we extend this norm to matrices $\operatorname{Mat}_{m\times m}(\mathbb{Y}_{c})$, again by taking the max of the norm of each coordinate. Note that this matrix norm is sub-multiplicative, so for $B_{1},B_{2}\in\operatorname{Mat}_{m\times m}(\mathbb{Y}_{c})$ we have
@@
The proof of the first extension of $\delta_{0}^{N}$ is given in [18, Prop 2.5.8] and relies on the fact that the map $\delta_{0}^{N}$ factors through $N/J^{d}N$ ($J$ is the maximal ideal generated by $(1\otimes a)-(a\otimes 1)$). For the second, we note that $M/\tau M$ is a $J$-torsion $L\otimes A$-module, so $\delta_{0}^{M}M\to M/\tau M\cong\operatorname{Hom}_{L}(\operatorname{Lie}_{\phi}(L),L)$ factors uniquely as
@@
We let $N\langle\sigma\rangle$ denote the subset of $n\in N\otimes_{L[\sigma]}L[[\sigma]]$ such that we have $(c_{1,j},\dots,c_{d,j})^{(j)}\to 0$ as $j\to\infty$ in the above expansion for $n$. Thus there exists an extension of the map $\delta_{1}^{N}:N\langle\sigma\rangle\to\mathbb{C}_{\infty}^{d}$ defined by applying $\delta_{1}^{N}$ to partial sums of (2.9), then taking the limit. For the $A$-motive $M$ and for fixed $\mathbf{z}\in L^{d}$, we define $M\langle\tau\rangle$ similarly, except subject to the condition that $(c_{1,j},\dots,c_{d,j})\mathbf{z}^{(j)}\to 0$. As above, there exists an extension of $\delta_{1,\mathbf{z}}^{M}$ to $M\langle\tau\rangle$.
@@ 3. The Exponential Motivic Pairing
The goal of this section is to define a pairing between the rings $(L\otimes_{\mathbb{F}_{q}}A)[\tau]$ and $(L\otimes_{\mathbb{F}_{q}}A)[\sigma]$, as well as a related pairing between the motives $M$ and $N$, which specializes to give the exponential function. This will allow us to give very explicit formulas for the coefficients of the exponential function for any abelian $\Xi$-finite Anderson $A$-module.
@@
We recall for the reader’s convenience that $\delta_{0}^{M}(g)\in L^{d}$ for any $g\in M_{(J)}$. Thus, in the above summand, $\delta_{0}^{M}(\tau^{-i}(x(g_{k})))^{\top}\mathbf{z}\in L$, and thus it makes sense to multiply this by $\sigma^{i}(y(h_{k}))\in N$. This is an element of $N$, so we can evaluate it under $\delta_{1}^{N}$, and taking the sum for $i\geq 0$ gives an element $\operatorname{Mat}_{d}(L)[[\mathbf{z}]]$. The question remains whether this sum converges to an element in $\mathbb{C}_{\infty}^{d}$. For now, we view the pairing as a formal sum in $\operatorname{Mat}_{d}(L)[[\mathbf{z}]]$, and we will address the convergence issues in Lemma 3.11 after establishing some of the properties of $F$.
@@
We thank the referee for the following helpful suggestion. It is possible to give a completely coordinate free definition of (3.2). Namely, for $m\in M$ and $n\in N$ and $\mathbf{z}\in\operatorname{Lie}_{E}(L)$ we set
@@
where we comment that this definition makes sense because we identify $M/\tau M\cong\operatorname{Hom}_{L}(\operatorname{Lie}_{E}(L),L)$ as in [18, 2.5.2], thus $\delta_{0}^{M}(\tau^{-i}m)(z)$ makes sense and is in $L$. Then, $N$ is an $L$-vector space, so $\delta_{0}^{M}(\tau^{-i}m)(z)\cdot\sigma^{i}n\in N$. Finally, $\delta_{1}^{N}\left(\delta_{0}^{M}(\tau^{-i}m)(z)\cdot\sigma^{i}n\right)\in N/(1-\sigma)N\cong E(L)$. Then, to recover Definition 3.1 we evaluate the above function $F$ at a chosen basis for $M$ and its dual basis for $N$ and take the trace.
We comment that, $\tau^{-i}m\in M_{(J)}$ always, with no restrictions on $M$. So with this construction, the $\Xi$-regularity condition of Definition 2.11 is not needed. Only once we choose bases for $M$ and $N$ as $\operatorname{Frac}(L\otimes_{\mathbb{F}_{q}}A)$-vector spaces, is the $\Xi$-regularity condition needed to evaluate this representation of $\tau^{-i}m$ at $\Xi$.
@@
For $\mathbf{z}\in L^{d}$, as formal series in $\operatorname{Mat}_{d}(L)[[\mathbf{z}]]$ we have
@@
Since $F(1,1;\mathbf{z})$ and $\operatorname{Exp}_{\phi}(\mathbf{z})$ agree as formal power series, and since we know $\operatorname{Exp}_{\phi}(\mathbf{z})$ is an entire function on $\mathbb{C}_{\infty}^{d}$ (see [1]), we conclude that $F(1,1;\mathbf{z})$ is also an entire function of $\mathbf{z}$. By Proposition 3.6 we see that $F(cx,dx;\mathbf{z})=cdF(x,y;\mathbf{z})$ for $c,d\in L$, that $F(ax,bx;\mathbf{z})=\phi_{b}\circ F(x,y;d[a]\mathbf{z})$ for $a,b\in A$. Further, recall that for $c\in\mathbb{C}_{\infty}$ and $h_{k}$ a basis element of Lemma 2.10, if $\delta_{1}^{N}(c\sigma^{i}(h_{k}))=w\in\mathbb{C}_{\infty}^{d}$, then by (2.4) we have that $\delta_{1}^{N}(c\sigma^{i+1}(h_{k}))=w^{(1)}$. From this, it follows that $F(\tau,1;\mathbf{z})=F(1,\sigma;\mathbf{z})=F(1,1;\mathbf{z})^{(1)}$. Thus, the convergence of $F(x,y;\mathbf{z})$ reduces to the convergence of $F(1,1;\mathbf{z})$, which is an entire function. Thus $\sum_{i=0}^{\infty}\sum_{k=1}^{d}\delta_{0}^{M}(\tau^{-i}(x(g_{k})))^{\top}\mathbf{z}\sigma^{i}(y(h_{k})))\in N\langle\sigma\rangle$.
@@
In this example we give details of many of the above constructions in the case of the $n$th tensor power of the Carlitz module for fixed $n>0$. We have $K=\mathbb{F}_{q}(t)$ and $A=\mathbb{F}_{q}[t]$ and we set $L=\mathbb{C}_{\infty}$. Let us identify $1\otimes A=\mathbb{F}_{q}[t]$ and $A\otimes 1=\mathbb{F}_{q}[\theta]$ for independent variables $t$ and $\theta$, thus obviating the need for the tensor product in this situation (thus $\mathbb{C}_{\infty}$ is written using the $\theta$ variable). We may realize the $t$-motive $M$ and the dual $t$-motive $N$ as $M=N=\mathbb{C}_{\infty}[t]$ (here $L\otimes_{\mathbb{F}_{q}}A=\mathbb{C}_{\infty}\otimes_{\mathbb{F}_{q}}A=\mathbb{C}_{\infty}[t]$) with
@@
for $z\in M_{K}$. Finally, Corollary 3.9 (here $d=1$) shows that the exponential coefficients $Q_{i}$ are given by
$Q_{i}=\delta_{0}^{M}(\tau^{-i}(1))=\left(\frac{1}{f^{(-1)}\cdots f^{(-i)}}\bigg{|}_{\Xi}\right)^{(i)}=\frac{1}{f\cdots f^{(i-1)}}\bigg{|}_{\Xi^{(i)}},$